Find four consecutive even integers such that the sum of twice the first, five times the second, and four times the third divided by the three times the fourth equals three.
I have tried this problem severall times. Some of the answers I got were, x= 0.2916, x- 70/111, x= -71/75,and x= 71/75. Please help and if possible include step-by-step instructions. Thank you!

Answer by mananth(16946)

let the numbers be x,x+2,x+4,x+6
..
2x+5(x+2)+4(x+4)/3(x+6)=3
2x+5x+10+4x+16=9(x+6)
11x+26= 9x+54
11x-9x=54-26
2x=28
x= 14
.....
14,16,18,20